OpenStudy (anonymous):

A typical laboratory centrifuge rotates at 4500 rpm. Test tubes have to be placed into a centrifuge very carefully because of the very large accelerations. (a) What is the acceleration at the end of a test tube that is 8 cm from the axis of rotation? m/s2 (b) For comparison, what is the magnitude of the acceleration a test tube would experience if dropped from a height of 1.2 m and stopped in a 1.0 ms long encounter with a hard floor? m/s2

OpenStudy (shamim):

a) Required formula is a=omega^2*r

OpenStudy (shamim):

Omega=2*pi*f

OpenStudy (shamim):

Frequency=f=4500rpm=4500/60hz

OpenStudy (caominhim):

A) \[A_{c} = \frac{ V^{2} }{r }\] B) \[K = \frac{ 1 }{2 }MV^{2}\]

OpenStudy (shamim):

r=8cm=0.08m a=?

OpenStudy (shamim):

R u able to do b)